OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. Answer x – 3y = –9 3x + y = 9 x – 3y = –8 3x – 3y = –8

OpenStudy (rulnick):

x-3y=-8

OpenStudy (accessdenied):

Notes for this problem: Slope: \[m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}} given: (x_1,y_1), (x_2,y_2) \] Point-slope form \[y - y_1 = m(x - x_1), given: (x_1,y_1)\] Once you have the equation of the line, you only need to move the x and y terms to one side and the constant terms to the opposite side (getting rid of any fractions through multiplication if necessary).